# Visual and auditory insights into multifrequency variability

### v2

@@ 1 Introduction
Data visualisation, often referred to as DataViz, serves as a valuable approach for identifying and discerning patterns or recurring trends within the behaviours of social or natural phenomena. It leverages its cognitive and explanatory capabilities through techniques such as time series, histograms, infographics, and Venn-Euler diagrams (Friendly, 2008). Consequently, graphical representation stands as a fundamental method for conducting research in the realms of science and study.
@@
DataViz stands as a important methodology for both scientific communication and research. Nonetheless, it is essential to consider a couple of factors. Firstly, it is important to recognise that visual communication is not accessible to individuals who are blind or have visual impairments (BVI) (Pérez-Montero, 2019). Secondly, relying solely on graphical representations might prove insufficient when dealing with comprehensive data analysis. Therefore, incorporating supplementary review methods becomes indispensable to enhance the research process.
@@
Taking this into account, data sonification can serve not only as a means of inclusivity in scientific communication but also as an additional tool for research (Sawe et al., 2020). By transforming data into sound, it offers an alternative mode of perception that complements visual representations, allowing for a deeper understanding of complex data.
As Kramer et al. (1999) defines it, data sonification as “the transformation of data relations into perceived relations in an acoustic signal for the purposes of facilitating communication or interpretation” highlights sonification as a valuable supplementary instrument for comprehending information. Unlike the sense of vision, hearing excels in perceiving information over time, as well as recognising patterns and momentary fluctuations. Therefore, sonification proves to be an excellent aid in enhancing knowledge understanding (Guttman et al., 2005).
@@
Another definition of data sonification, provided by Hermann (2008), characterises it as a technique that takes data as its input and produces sound signals, possibly in response to optional additional stimulation or triggering. This broad definition encompasses any process that uses data to create auditory outputs and can be referred to as sonification, if and only if the following conditions are satisfied:
@@
  • The system can intentionally be used with different data and also be used in repetition with the same data.
@@
Following Hermann (2008) and Vogt (2008), the “sonification is the data-dependent generation of sound, if the transformation is systematic, objective and reproducible, so that it can be used as scientific method.”
Numerous instances demonstrate the effectiveness of data sonification as a valuable science communication tool. Notable examples include its application in depicting salmon migration patterns (Hegg et al., 2018) and visualising fluctuations in brainwave activity (Parvizi et al., 2018). The Geiger counter represents yet another compelling case in point.
@@
In the realm of astronomy, early instances of data sonification trace to Jansky (1933), when detected radio waves emanating from the centre of the Milky Way were analysed by telephone communication noise. Three decades later, Penzias & Wilson (1965) identified the interference in their observations as originating from cosmic microwave background radiation, marking a significant discovery in the field (Kellermann et al., 2020).
@@
Another example is that of Morgan et al. (1997) who took the data of the X-ray emission of the black hole GRS 1915+105 and translated it into audio signals, allowing us to “hear” the accretion disk from it Masetti (2013). Also Chandra (2003) has sonified data and released “Hears” of a black hole for the first time.
@@
Abbott et al. (2016) documented the simultaneous detection of a brief signal by two gravitational wave detectors, which resulted from the fusion of two black holes. The researchers converted the data from the instant of the collision, which they dubbed a “chirp.” This term was chosen because certain events producing gravitational waves exhibit auditory similarities to the chirping sounds of birds (LIGO Collaboration, 2016, see https://youtu.be/QyDcTbR-kEA).
@@
Furthermore, Berti (2016) published an article entitled The First Sounds of Merging Black Holes, shedding light on the initial auditory cues of black hole mergers. In recent times, the National Aeronautics and Space Administration (NASA) has unveiled sonification imagery derived from optical data captured by the Hubble Space Telescope (NASA, 2022) and X-ray data observed by the Chandra X-ray Observatory (CXO) (Chandra, 2020).
In the context of communicating science to the BVI community, especially in the realm of astrophysics and black holes (BHs) this work sonifies data from nine blazars: Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153.
[…]
@@
With this background, in Section 2 we show the way to obtain the multifrequency data (radio, optical, X-ray, and $\gamma$-ray). Section 3 shows the sonification procedure: pre-production, production, and post-production. The discussion is shown in Section 4 with which we conclude our study.
@@ 2.1 Radio
The radio datasets for this work were obtained from two data bases: the University of Michigan Radio Astronomy Observatory (UMRAO) [^1] https://dept.astro.lsa.umich.edu/datasets/umrao.php, and the Astrogeo Very-long-baseline interferometry (Astrogeo-VLBI).
The 26-meter paraboloid UMRAO monitoring in the bands: 4.8, 8, and 14.5 GHz (Aller et al., 1985). It is focused on the study of total flux density and linear polarisation of active extragalactic objects [^2] https://lsa.umich.edu/astro/facilities/data-sets.html.
Radio Fundamental Catalog (RFC), of the Astro-geo Center, maintained by L. Petrov333 http://astrogeo.org/rfc provides precise positions with milli-arcseconds accuracies, maps, and estimates of correlated flux densities at baselines 1000–8000 km of thousands of compact radio sources.
@@ 2.2 Optical
The optical AAVSO is a public database founded in 1910 and offers observations of variable stars collected and archived for world-wide access in collaboration with amateur and professional astronomers. It is an international organisation of long-term variable star observers who participate in scientific discovery through variable star astronomy.
@@ 2.3 X-rays
For the X-rays light curve, we use the Swift database with energies in the range $0.3-10\,\text{keV}$. It contains the X-Ray Telescope (XRT) with two important characteristics that makes it important for observations: low background and a constant point spread function across the field of view (Moretti et al., 2005).
@@ 2.4 Gamma-rays
The public database of $\gamma$-rays used was of The Fermi Large Area Telescope (LAT). This instrument observes the sky for $\sim$ 3 hours and covers $\gtrsim$ 20% of it for each observation. It has a crystal calorimeter for energy measurement, an anti-coincidence detector to distinguish the background of charged particles and high-resolution converter tracker for direction measurement of the incident gamma rays (Atwood et al., 2009). The $\gamma$-rays fluxes collected are in the range $100\,\text{MeV}-300\text{GeV}$.
@@ 3 Sonification
Sonifying the data for each of the blazar frequencies was carried out in three stages: pre-production, production and post-production.
@@ 3.1 Pre-production
Sonification pre-production basically focuses on data pre-processing. In the case of the radio and optical bands, the data was processed in such a way that any anomaly, disturbance or irrelevance of the data could be detected, eliminated, corrected or transformed. For this task the AWK pattern scanning and processing language was used.
In the case of X-rays, it was necessary to reduce the data and for this, the HEASoft [^6] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/ software was used. The data processing corresponding to the photons, the position and the orientation of the satellite is available on Flexible Image Transport System format files (FITS).
For the pre-production of sonification in the $\gamma$-rays, the data reduction was possible with the Fermitools [^7] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ software. For this, it was necessary to know the main relevant parameters of the blazars such as: R.A., Dec., Region of Interest (ROI) and start and end time of the data processing (see e.g Magallanes-Guijón, 2020). The process of this reduction followed the procedure detailed by Cabrera et al. (2013).
@@
3.1.1 Light curves
Each blazar selected had the following specifications:
Band Records Interim Total
  [y, m, d] [d]
Optical $11,849$ $23,2,17$ $8441$
X-rays $28,000$ $12,7,9$ $4607$
$\gamma$-rays $4,199$ $11,7,9$ $4239$
Table 1: The table shows each electromagnetic band studied for Mrk 501: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
@@ 3.1.1 Light curves
Band Records Interim Total
 [y, m, d] [d]
Optical $1255$ $11,4,0$ $4141$
X-rays $11150$ $12,5,20$ $4554$
$\gamma$-rays $5146$ $14,2,18$ $5192$
Table 2: The table shows each electromagnetic band studied for Mrk 1501: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least one measurement is taken [d].
@@
Band Records Interim Total
 [y, m, d] [d]
Radio $413$ $12,5,25$ $4560$
Optical $10117$ $52,2,27$ $19082$
X-rays $100418$ $14,1,8$ $5152$
$\gamma$-rays $4892$ $13,7,7$ $4967$
Table 3: The table shows each electromagnetic band studied for BL Lacertae: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
@@
[…]
Band Records Interim Total
 [y, m, d] [d]
Optical $196$ $22,9,17$ $8055$
X-rays $61146$ $9,2,5$ $3353$
$\gamma$-rays $4223$ $11,8,1$ $4262$
Table 4: The table shows each electromagnetic band studied for AO 0235+164: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
Figure 4: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar AO 0235+164. The sonification is available in %****␣magallanes-mendoza.tex␣Line␣575␣****https://www.guijongustavo.org/datasonification/ao0335/playlist.html
Band Records Interim Total
 [y, m, d] [d]
Radio $101$ $23,4,22$ $8545$
Optical $8470$ $28,0,9$ $10236$
X-rays $3163$ $14,2,7$ $5181$
$\gamma$-rays $4223$ $11,8,1$ $4262$
Table 5: The table shows each electromagnetic band studied for 3C 66A: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
@@
Band Records Interim Total
 [y, m, d] [d]
Optical $76$ $17,5,22$ $6383$
X-rays $16111$ $15,2,8$ $5548$
$\gamma$-rays $4196$ $11,8,1$ $4262$
Table 6: The table shows each electromagnetic band studied for 0J 049: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
@@
(7) PKS J2134-0153 has a redshift $z$ = 1.285, corresponding to a distance $\sim$ 5 Gpc (Truebenbach & Darling, 2017) with R. A. = 21h 34m 10.3095s and Dec. = $-01^{\circ}$ 53’ 17.238” (Fomalont et al., 2000).
Band Records Interim Total
 [y, m, d] [d]
Optical $73$ $0,10,10$ $314$
X-rays $4123$ $6,0,6$ $2197$
$\gamma$-rays $4223$ $11,8,1$ $4262$
Table 7: The table shows each electromagnetic band studied for PKS J2134-0153: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
@@
Band Records Interim Total
 [y, m, d] [d]
Optical $18167$ $35,9,20$ $13079$
X-rays $160346$ $15,1,0$ $5510$
$\gamma$-rays $4129$ $11,9,3$ $4294$
Table 8: The table shows each electromagnetic band studied for Mrk 421: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].
@@
Band Records Interim Total
 [y, m, d] [d]
Optical $13754$ $32,8,6$ $11939$
X-rays $9464$ $11,5,16$ $4185$
$\gamma$-rays $4410$ $12,2,2$ $4446$
Table 9: The table shows each electromagnetic band studied for OJ 287: the number of records, the interim in years, months, and days [y, m, d], and the total days for which at least a measurement is obtained [d].

### v3

@@ 1 Introduction
Data visualisation, often referred to as DataViz, serves as a valuable approach for identifying and discerning patterns or recurring trends within the behaviours of social or natural phenomena. It leverages its cognitive and explanatory capabilities through techniques such as time series, histograms, infographics, and Venn-Euler diagrams (Friendly, 2008). Consequently, graphical representation stands as a fundamental method for conducting research in the realms of science and study. Moreover, its ubiquity has made it not only an analytical tool but also the default and almost universal medium for communicating information across scientific, educational, and even everyday contexts, where visualisation often constitutes the first and most natural choice for conveying knowledge (Harris, 1999).
@@
DataViz stands as an important methodology for both scientific communication and research. Nonetheless, it is essential to consider a couple of factors. Firstly, it is important to recognise that visual communication is not accessible to individuals who are blind or have visual impairments (BVI) (Pérez-Montero, 2019). Secondly, relying solely on graphical representations might prove insufficient when dealing with comprehensive data analysis, which involves exploring large, multidimensional datasets, detecting subtle correlations, or interpreting complex temporal and spatial patterns. Therefore, incorporating supplementary review methods—such as statistical modelling, sonification, and interactive data exploration—becomes indispensable to ensure a more complete and robust understanding of the data (Hearst, 2011).
@@
Data sonification can serve not only as a means of inclusivity in scientific communication but also as an additional tool for research (Sawe et al., 2020). By transforming data into sound, it offers an alternative mode of perception that complements visual representations, allowing for a deeper understanding of complex data.
Building on the limitations of purely visual approaches, data sonification and sound visualization have emerged as complementary methodologies for exploring and analyzing complex datasets. Notably, Kaper et al. (2000) have demonstrated how digital sound synthesis, via tools such as DIASS (Digital Instrument for Additive Sound Synthesis), combined with immersive sound visualization in virtual-reality environments like M4CAVE, can facilitate the identification of patterns, correlations, and anomalies in multidimensional scientific data. Their work illustrates that auditory representations not only enhance inclusivity for blind or visually impaired (BVI) individuals but also provide all researchers with an additional perceptual channel, allowing subtle temporal and structural features in the data to be perceived more effectively.
In this context, Kramer et al. (1999) define data sonification as “the transformation of data relations into perceived relations in an acoustic signal for the purposes of facilitating communication or interpretation.” This definition positions sonification as a supplementary tool that can aid in understanding complex datasets. Research indicates that the human auditory system is particularly effective at perceiving information distributed over time and detecting temporal patterns, often outperforming vision in these tasks (McAdams & Bigand, 1993). Consequently, sonification proves to be an effective method for enhancing comprehension of temporal and sequential information, revealing trends, fluctuations, and correlations in datasets that might be less apparent in visual representations alone  (Guttman et al., 2005).
@@
Another definition of data sonification, provided by Hermann (2008), characterises it as a technique that takes data as its input and produces sound signals. In some cases, the generation of these sounds may be influenced by additional stimuli or triggers, which are optional and not strictly necessary for the process. This broad definition considers as sonification any process that uses data to create auditory output if the following conditions are satisfied:
@@
  • The system is designed to be reusable and generalizable, meaning it can be applied to different datasets or used iteratively to verify consistency with the same dataset.
@@
Following Hermann (2008) and Vogt (2008), this paper adopts the definition of sonification as the data-dependent generation of sound, provided that the transformation is systematic, objective, and reproducible, so that it can be used as a scientific method. This definition is chosen because it emphasizes the methodological rigor required for sonification to be considered a reliable tool for scientific analysis, distinguishing it from more artistic or exploratory uses of sound.
Numerous instances demonstrate the effectiveness of data sonification as a valuable science communication tool. For the purposes of this paper, science communication refers to the process of conveying scientific information and insights to diverse audiences—researchers, students, and the general public—in a manner that is both understandable and engaging. Notable examples include its application in depicting salmon migration patterns  (Hegg et al., 2018) and visualising fluctuations in brainwave activity (Parvizi et al., 2018).Another compelling case is the Geiger counter, a device used to detect and measure ionizing radiation. In traditional use, the counter produces audible clicks corresponding to radiation events, effectively sonifying the underlying radioactive activity. This auditory representation allows listeners to perceive variations in radiation intensity over time, providing both an immediate, intuitive understanding of the data and a practical tool for real-time monitoring.
@@
In the realm of astronomy, early connections between sound and scientific data can be traced to Jansky (1933), who first detected radio waves emanating from the centre of the Milky Way and identified them through the auditory noise in telephone communications. While Penzias & Wilson (1965) three decades later detected the cosmic microwave background radiation using radio instrumentation, their work exemplifies how radio data—initially experienced as noise—would later inspire the development of auditory and sonification approaches in astronomy (Kellermann et al., 2020).
@@
Another example is that of Morgan et al. (1997) who took the data of the X-ray emission of the black hole GRS 1915+105 and translated it into audio signals, allowing us to “hear” the accretion disk from it Masetti (2013). Also Chandra (2003) has sonified data and released “Chandra ‘Hears’ of a black hole for the first time”.
@@
Abbott et al. (2016) documented the simultaneous detection of a brief signal by two gravitational wave detectors, which resulted from the fusion of two black holes. The researchers converted the data from the instant of the collision, which they dubbed a “chirp,” a term chosen because the sonified signal exhibits auditory similarities to the chirping sounds of birds (LIGO Collaboration, 2016, see https://youtu.be/QyDcTbR-kEA), even though this sound arises solely from the sonification method and not from any actual audible event. Furthermore, Berti (2016) published an article entitled “The First Sounds of Merging Black Holes”, shedding light on the initial auditory cues of black hole mergers.
@@
In recent times, the National Aeronautics and Space Administration (NASA) has unveiled sonification of imagery derived from optical data captured by the Hubble Space Telescope (NASA, 2022) and X-ray data observed by the Chandra X-ray Observatory (CXO) (Chandra, 2020).
Recent advances in astronomical sonification have demonstrated the potential of combining auditory and visual data representations. For instance, Tucker Brown et al. (2022a) assessed Astronify (Brasseur et al., 2023), a prototype tool integrated within the MAST archive that enables sonification of evenly sampled light curves. Their study evaluated the effectiveness of sonification for signal detection at various signal-to-noise ratios, showing that auditory representations can aid in identifying high-SNR transits and complement traditional visual inspection. More recently, Huppenkothen et al. (2023) presented the Sonified Hertzsprung–Russell Diagram, which transforms stellar time series into sound, encoding physically meaningful features that preserve astrophysical differences between stars through both visual and auditory media. Our work extends these approaches by applying sonification to multiwavelength light curves of blazars, integrating light curves, waveforms, and spectrograms to explore temporal evolution across distinct energy regimes (optical, X-ray, and $\Gamma$-ray). This multifrequency perspective highlights variability patterns in active galactic nuclei and demonstrates how sonification can support both scientific interpretation and public engagement in high-energy astrophysics.
In the context of communicating science to the BVI community, especially in the realm of astrophysics and black holes (BHs) this work sonifies data from nine blazars: Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153. These objects have been reported in the literature as exhibiting quasi-periodic variability in different electromagnetic bands. By sonifying their light curves, our aim is to explore whether auditory perception can aid in identifying, characterising, and communicating such periodicities and variability patterns. Further details are given in Appendices A and B.
The selection of blazars analyzed in this study was based on three main criteria: (i) the availability of long-term, sampled light curves across multiple wavelengths (radio, optical, X-ray, and gamma-ray), (ii) documented evidence of significant variability on different timescales, and (iii) their inclusion in public monitoring programs that ensure data accessibility and reproducibility. The chosen sources—such as Mrk 421, Mrk 501, and 3C 273—are among the most extensively observed and best-characterized blazars in the literature, known for their pronounced flux variations and strong multi-band correlations (Abdo et al., 2011). These properties make them ideal test cases for exploring how sonification can enhance the perception and interpretation of temporal patterns in multifrequency datasets.
The ensemble of data employed in this study was selected to encompass a broad range of energy bands (radio, optical, X-ray, and $\gamma$-ray) in order to capture the multifrequency variability characteristic of blazars. While the temporal coverage is not fully simultaneous across all bands, the datasets from the Swift and Fermi observatories are largely contemporaneous, as both missions were launched in 2008, providing overlapping X-ray and gamma-ray observations over extended periods. This temporal alignment allows for meaningful cross-band comparisons in several sources. For other frequency ranges, such as optical and radio, the data were chosen to complement these high-energy observations, offering insights into long-term variability trends. From a sonification perspective, these differences in time sampling and frequency response enrich the auditory experience, allowing users to perceive asynchronous variability patterns, phase shifts, and multiwavelength correlations through sound.
This sonification serves a dual goal: it enables BVI audiences to directly engage with astrophysical data, perceiving temporal patterns and variability that are typically conveyed visually, while also providing astronomers with an analytical tool to detect trends and anomalies in light curves through auditory perception, thereby complementing conventional visual analysis methods.
Although this study does not aim to definitively demonstrate periodicities, the auditory representation provides a means to detect potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects. By demonstrating how sonification can complement traditional visualizations, this work establishes a methodological foundation for future research, including systematic searches for periodicities, transient events, and multiwavelength correlations in astronomical data. This multimodal approach enables a more inclusive and perceptually rich exploration of the variability of blazars for both sighted and visually impaired audiences.
Blazars are extragalactic sources of intense luminosity, with energy outputs ranging from $10^{41}$ to $10^{47}$ ergs s-1 (Blandford et al., 2019), whose relativistic jets are oriented close to the observer’s line of sight (Ulrich et al., 1997). This geometric alignment results in strong Doppler boosting, which amplifies the observed emission and produces rapid variability across the electromagnetic (EM) spectrum. Their radiation is predominantly non-thermal and arises from two main mechanisms: synchrotron emission, which dominates the radio-to-optical range, and inverse Compton scattering, responsible for X-ray and gamma-ray photons (Padovani et al., 2012).Their pronounced variability over a wide range of timescales makes them particularly suitable for this study, where they serve as test cases to assess the potential of sonification both as an analytical tool and as a medium for science communication.
The variability observed at different wavelengths reflects physical changes in distinct regions of the jet — for instance, fluctuations in magnetic field strength, particle acceleration, or shock propagation — leading to complex temporal structures that can range from minutes to years. Because these variations encode valuable information about the underlying astrophysical processes, blazars represent ideal candidates for data sonification. Transforming their multi-band light curves into sound provides an alternative perceptual channel for detecting periodicities, correlations, or transient events that might be less evident through visual inspection alone.
@@
With this background, in Section 2 we describe how we obtained the multifrequency data (radio, optical, X-ray, and $\gamma$-ray). Section 3 shows the sonification procedure: pre-production, production, and post-production. The discussion is shown in Section 4 with which we conclude our study.
@@ 2.1 Radio
The radio datasets for this work were obtained from two data bases: the University of Michigan Radio Astronomy Observatory (UMRAO) [^1] https://dept.astro.lsa.umich.edu/datasets/umrao.php, and the Astrogeo Very-long-baseline interferometry (Astrogeo-VLBI) database.
From UMRAO, we obtained long-term monitoring light curves of total flux density and linear polarization at 4.8, 8.0, and 14.5 GHz measured with the 26 m radio telescope (Aller et al., 1985). For the sonification, data points from the three frequencies were not averaged or combined interchangeably; instead, they were arranged in temporal order while preserving their associated frequency tags. This approach allowed us to create a continuous auditory sequence reflecting the multifrequency monitoring cadence, where each tone corresponds to an observation at a specific band. The selected frequencies are those used in UMRAO’s long-term variability program, optimized for compact radio sources. This structure made it possible to represent the variability across frequencies and identify temporal patterns or irregular sampling intervals through sound. [^2] https://lsa.umich.edu/astro/facilities/data-sets.html The Astrogeo-VLBI data, corresponding to a frequency of 8.4 GHz (Petrov & Kovalev, 2025), provides complementary spatial information through very long baseline interferometry.
The Radio Fundamental Catalog (RFC), of the Astro-geo Center, maintained by L. Petrov [^3] http://astrogeo.org/rfc provides precise positions with milli-arcseconds accuracies, maps, and estimates of correlated flux densities for thousands of compact radio sources. The mentioned baselines of 1000–8000 km mentioned in the catalog (RFC) correspond to the distances between the radio telescopes in the very long baseline interferometry (VLBI) network. Longer baselines provide higher angular resolution, allowing for finer spatial details of the sources. These datasets also include light curves provided in tabular format (flux density vs. time), which are suitable for variability analysis.
@@ 2.2 Optical
From AAVSO, we obtained publicly available photometric light curves of blazars in different optical filters, contributed by both amateur and professional astronomers. The data are distributed as tab-delimited American Standard Code for Information Interchange (ASCII) tables containing Julian Date, magnitude, filter, and error. The optical AAVSO is a public database founded in 1910 and offers observations of variable stars collected and archived for world-wide access in collaboration with amateur and professional astronomers. It is an international organisation of long-term variable star observers who participate in scientific discovery through variable star astronomy.
@@ 2.3 X-rays
For the X-ray light curves, we used data from the Swift database covering energies in the range 0.3–10 keV. The X-Ray Telescope (XRT) has two key characteristics that are crucial for this analysis: a low background and a nearly constant point spread function (PSF) across the field of view. A constant PSF ensures that the spatial distribution of X-ray photons from the source is measured uniformly across the detector, reducing systematic effects when constructing light curves and enabling more reliable detection of flux variations (Moretti et al., 2005). The selected blazars are relatively X-ray bright, providing sufficient photon counts to generate meaningful temporal variations for both visualization and sonification. The Swift database provides background-subtracted light curves in Flexible Image Transport System (FITS) and ASCII table formats, including time, count rate, and errors. These products were directly used in our analysis with additional reduction.
@@ 2.4 Gamma-rays
For the gamma-ray regime, we used data from the Fermi Large Area Telescope (LAT) in the 100 MeV–300 GeV range. Specifically, we used binned light curves in FITS files from the public Fermi-LAT Light Curve Repository, which provides fluxes integrated over pre-defined time bins (typically weekly) in tabular format (flux vs. time with associated errors). These products are already processed through the LAT standard analysis pipeline FERMITOOLS [^6] https://fermi.gsfc.nasa.gov/ssc/data/analysis/software/ and suitable for variability studies. This instrument observes the sky for $\sim$ 3 hours and covers $\gtrsim$ 20% of it for each observation. It has a crystal calorimeter for energy measurement, an anti-coincidence detector to distinguish the background of charged particles and high-resolution converter tracker for direction measurement of the incident gamma-rays (Atwood et al., 2009).
@@ 3 Sonification
Sonifying the data for each of the blazar frequencies was carried out in three stages: pre-production, production and post-production detailed in the next section.
@@ 3.1 Pre-production
  • Eliminated if the value was clearly spurious (e.g., magnitude far outside the instrument’s dynamic range, corrupted entries).
  • Corrected if a small inconsistency could be addressed (e.g., date formatting or duplicated entries consolidated).
  • Transformed if a re-scaling or homogenization was required (e.g., converting magnitudes to flux densities, or aligning different photometric systems).
This preprocessing was performed using the AWK pattern scanning and processing language, which allows filtering, restructuring, and formatting the data into consistent light curve tables.
For the X-ray data, a full reduction was necessary. Raw event files were processed with the HEASoft [^7] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/ package, which applies the standard calibration pipelines (e.g., bias subtraction, exposure corrections, effective area, and deadtime corrections). The photon event information, together with spacecraft attitude (position and orientation), is stored in FITS files, which serve as the standard format for astronomical data exchange and allow reproducibility of the analysis.
For the $\gamma$-ray data, reduction was performed using the Fermitools software suite. The procedure required specifying the relevant astrophysical parameters of each blazar—Right Ascension (R.A.), Declination (Dec.), Region of Interest (ROI), and the start and end times of the data processing—following the standard pipeline described by Cabrera et al. (2013) and Magallanes-Guijón (2020).
[…]
@@
Figure 4: From top to bottom, left to right, the Figure shows panels of optical, X-ray and $\gamma$-rays light curves, waveforms of the sonification as a function of time, and spectrograms of the blazar AO 0235+164. The sonification is available in https://www.guijongustavo.org/datasonification/ao0335/playlist.html